100g of a mixture of NaOH and Na2SO4 is neutralised by 100ml of 0.5M H2SO4. What amount of Na2SO4 is present in the mixture.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Let moles of NaOH and Na2SO4 be x and y respectively.
moles x v.f = constant.
x * 1 + y*2 = 100*0.5 * 2
x + 2y = 100..................(1)
x = z/40  and  y = (100-z)/142 ….............(2)
put eqns. (2) in (1). and get the variable z solved. hence.